Find the function of the form y=loga^x whose graph is given. I was given the point (1/2, -1) Thank you for your time.
Answer by rothauserc(4718)

using point (1/2, -1), x = 0.5 and y = -1, therefore
-1 = log base a (0.5)
using definition of logarithms, we get
0.5 = a^-1
1/a = 0.5
multiply both sides of = by a (note a is not = 0)
0.5a = 1
a = 1/0.5 = 2
therefore the log equation is
y = log base 2 (x)
check answer
-1 = log base 2 (0.5) = -1
our answer checks